OpenStudy (anonymous):

integrate (x^3+x)/x^2 from 1 to 3

zepdrix (zepdrix):

\[\large \int\limits_1^3 \frac{x^3+x}{x^2}dx\]We can split this into two fractions,\[\large \int\limits\limits_1^3 \frac{x^3}{x^2}+\frac{x}{x^2}dx\]Then simply divide each term to simplify, and the apply the `Power Rule for Integration`. Understand how to proceed? :)

OpenStudy (anonymous):

no, unfortunately i don't know how to integrate. do i sub in 1, 2 and 3 for each x?

zepdrix (zepdrix):

After you've simplified you should have,\[\large \int\limits\limits\limits_1^3 x+\frac{1}{x}dx\] We'll integrate these separately, the same way we would do if we had to take a derivative with addition between the terms. The steps will be a little different though.

zepdrix (zepdrix):

Remember the `Power Rule for Derivatives`? It involves 2 steps: ~Multiply by the power. ~Decrease the power by one. The `Power Rule for Integration` will be the opposite, and in the opposite order. ~Raise the power by 1. ~Divide by the power.

zepdrix (zepdrix):

Here's a quick example,\[\large \int\limits x^3\;dx \qquad =\qquad \frac{x^{(3+1)}}{3+1} \qquad = \qquad \frac{1}{4}x^4\]

zepdrix (zepdrix):

So in your problem, that first term is \(x^1\). Try applying this rule to it.

OpenStudy (anonymous):

so would it look like this? \[\int\limits_{1}^{3}x+0 dx\]

zepdrix (zepdrix):

We can't apply the `Power Rule for Integration` to that second term. It's a special power that we're not allowed to apply the rule to. Only apply it to the first term. Did you try? :o

OpenStudy (anonymous):

yes. does that mean that it would be \[1/2\int\limits_{1}^{3}x^2\]

zepdrix (zepdrix):

Yes good c: Now for the second term, we'll need to think back to an important derivative.\[\large \left(\ln x\right)'=\frac{1}{x}\]The `derivative` of ln x gives us this. So if we want the `Anti-derivative` of 1/x, what will we get? :)

OpenStudy (anonymous):

is it lnx? lol

zepdrix (zepdrix):

yessss c:

OpenStudy (anonymous):

great! ok now what do i do with the 3 and 1? do i sub in now?

zepdrix (zepdrix):

Remember when you would take derivatives, the derivative operator, or the prime, would disappear as a part of the process of finding the derivative. Same thing happens with integration, the Swirly S bar, and the dx will disappear as a part of the integration process. We'll show that we still need to plug in our limits by drawing a bar behind our anti-derivative that we've found.\[\large \int\limits\limits\limits\limits_1^3 x+\frac{1}{x}dx \qquad = \qquad \frac{1}{2}x^2+\ln x \;|_1^3\]

zepdrix (zepdrix):

So we'll plug in the 3, and subtract from that, the function with 1 plugged in.\[\large \frac{1}{2}x^2+\ln x \;|_1^3 =\] \[\huge \left(\frac{1}{2}\color{royalblue}{3}^2+\ln \color{royalblue}{3}\right)-\left(\frac{1}{2}\color{royalblue}{1}^2+\ln \color{royalblue}{1}\right)\]

OpenStudy (anonymous):

so is the answer\[4+\ln 2\]

zepdrix (zepdrix):

\[\large \ln1=0\]

OpenStudy (anonymous):

oh right! so the answer is just 4?

zepdrix (zepdrix):

Where did your ln3 go? :D \[\large 4+\ln3\]

OpenStudy (anonymous):

oh ok. i understand more now. thanks so much!

zepdrix (zepdrix):

yay c: